what is the vertex form for:y=1/3x^2+2x+3 ?
Answer by Onmyoji.S(32)

1/3x²+2x+3=y
3(1/3x²-2x+3)=3y
x²-6x+9=3y
(x-3)(x-3)=3y
(x-3)²=3y
(x-3)²/3=y